4. Find the square roots of the following numbers by the Prime Factorisation Method.   iv. 4096

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student



4096 = 2×2×2×2×2×2×2×2×2×2×2×2
⇒ 4096 = (2×2)×(2×2)×(2×2)×(2×2)×(2×2)×(2×2) ⇒ 4096 = (2×2×2×2×2×2)×(2×2×2×2×2×2)
⇒ 4096 = (2×2×2×2×2×2)2
⇒ √4096 = 2×2×2 ×2×2×2 = 64